The Department of Computer Science and Engineering

University of Nevada, Reno

cordially invites you to a Master’s colloquium

A Unified Approach for Cross-Platform Software Development

A thesis submitted in partial fulfillment of the requirements for the degree of Master of Science with a major in Computer Science.

by

Jeffery Alan Stuart

Abstract: Cross-platform software development is a complex and challenging activity. Frequently, developers have to create portions of code that use platform-specific data types and functions. This has led to two largely adopted practices: either making extensive use of the preprocessor, or splitting the software package into several branches, one for each target platform. Both practices have their drawbacks. To tackle the issues of cross-platform development, this thesis proposes two programming solutions referred to as cores and routers. By using them, the need for advanced preprocessing and separate development branches is virtually eliminated. The conceptual solutions are described and examples of application are presented.

1:00 pm, Wednesday, June 15, 2005

Access Grid Node, Scrugham Engineering and Mines (SEM) room 201

for more information contact Dr. Fred Harris @ 784-6571 (fredh@cs.unr.edu)